helga invested $1000 at 5% interest rate, compounded anually. how much interest had helga earned by the end of the sixth year?
Answer by rfer(10417)

A=P(1+.05)^6
A=1000(1.05)^6
A=1000(1.34)
A=1340.09
i=1340.09-1000=$340.09